# Clustering in Networks with Time-varying Nodal Attributes

### v1

@@ Abstract
This manuscript studies nodal clustering in graphs having a time series at each node. The framework includes priors for low-dimensional representations and a decoder that bridges the latent representations and time series. The structural and temporal patterns are fused into representations that facilitate clustering, addressing the limitation that the evolution of nodal attributes is often overlooked. Parameters are learned via maximum approximate likelihood, with a graph-fused LASSO regularization imposed on prior parameters. The optimization problem is solved via alternating direction method of multipliers; Langevin dynamics are employed for posterior inference. Simulation studies on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a book word co-occurrence network demonstrate the effectiveness of the proposed method.
Keywords: ADMM, LASSO, Latent Space Models, Representation Learning, Time Series.
@@ 1 Introduction
Networks are used to describe relational phenomena, where entities are denoted by nodes and relations are delineated by edges between nodes. Such structures naturally arise in various domains, including social interactions (snijders2011statistical), neurology (yang2022functional), and finance (jackson2021systemic). A fundamental task in network analysis involves clustering nodes into disjoint sets, where node within a set are more connected than nodes in disjoint sets. Such clusters often provide structural insights into the network (see shai2017case for a collection of case studies). The majority of clustering methods focus on structural patterns (abbe2018community). However, many networks possess additional “nodal attributes”. This paper shows how to take advantage of nodal time series, providing clustering methods that account for both network structure and attribute dynamics. Section 4 uses the methods to identify climate regions in California by county. Here, both spatial proximities and temperatures of the counties are used to develop a realistic clustering.
@@
Clustering nodes of networks with nodal attributes relies on effectively combining graphical structures with the nodal information. Methods have been previously proposed for clustering attributed networks. binkiewicz2017casc perform covariate-assisted spectral clustering by augmenting the Laplacian of the graph with pairwise similarities of the nodal covariates. shen2024bayesian propose a Bayesian stochastic block model where nodal information is incorporated into prior distributions via a covariate-dependent random partition prior. hu2024nac leverage individual and neighboring covariates with a node-specific weight matrix to improve clustering. Furthermore, the graph-fused LASSO regularizations in chen2023more and Padilla2023graphon have shown promising results, producing smoothed signals between neighboring nodes while preserving boundaries between regions having distinct signals. Following hallac2015network and yu2025graph, our framework also adapts graph-fused LASSO regularization to achieve similar filtering behavior.
@@
In many applications, nodal attributes progress in time, while the network structure remains fixed. Nodal time series can provide valuable clustering information, especially in cases with a non-informative graph. Some methods for clustering graphs nodes overlook the evolution of nodal features, eschewing an approach that merges temporal dynamics with structural patterns. A natural solution introduces a latent variable at each node. By compressing each nodal series into a low-dimensional latent representation, temporal dynamics can be summarized and integrated with structural patterns. In our approach, latent representations are managed by node-specific priors and empirical Bayes techniques. Simultaneously, the latent space is regularized by a graph-fused LASSO penalty to incorporate network information. The resulting low-dimensional representations preserve both neighbor information and attribute dynamics, promoting effective clustering.
@@
Latent space models bridge low-dimensional representations with the observed nodal series. handcock2007model model nodes in a Euclidean space by assuming that edge probabilities are greater for node pairs that are closer in the latent space. zhu2023disentangling construct a latent space model with both attractive and repulsive nodes, thereby disentangling network polarization dynamics. In contrast, our framework models the nodal series, conditioning on its low-dimensional representations. Unlike a variational auto-encoder (kingma2013auto), which encourages the approximate posterior to align with a zero-mean Gaussian prior, we focus on learning Gaussian prior means to facilitate clustering. With graph-fused LASSO regularization imposed on the differences between prior parameters over the edges, nodes having similar patterns are encouraged to have similar prior parameters. Consequently, a $k$-means algorithm that clusters via centroids is well-suited.
@@
  • An alternating direction method of multipliers (ADMM) procedure is developed to solve our associated optimization problem; parameters are learned via posterior inference. Through simulations on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a word co-occurrence network, the effectiveness of the proposed framework is demonstrated.
The rest of the manuscript is organized as follows. Section 2 describes our proposed framework. Section 3 presents an objective function with a graph-fused LASSO regularization and develops the ADMM procedure to solve the optimization problem. Section 4 illustrates the proposed method through simulations and real-data applications. Section 5 discusses limitations and ideas for future work.
@@ 2 Decoder-only Latent Space Models
We consider an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the node set and $\mathcal{E}$ is the edge set. For each node $i\in\mathcal{V}$, let $\{Y_{t,i}\}_{t=1}^{n}$ denote a sequence of univariate random variables (the node $i$ time series). The node $i$ series is represented as $\bm{Y}_{i}\coloneqq(Y_{1,i},Y_{2,i},\dots,Y_{n,i})^{\top}\in\mathbb{R}^{n}$. For each $\bm{Y}_{i}\in\mathbb{R}^{n}$, we assume existence of a latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ such that $\bm{Y}_{i}$ is generated from $\bm{Z}_{i}$ with the following time series “decoder”:
@@
where $\bm{h}_{\bm{\phi}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ is a neural network parameterized by $\bm{\phi}$. Here, $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ is the conditional probability density function of $\bm{Y}_{i}$ given $\bm{Z}_{i}$ (which depends on the neural network parameters in $\bm{\phi}$). We impose a prior distribution on the latent $\bm{Z}_{i}$ via
@@
The prior parameter, $\bm{\mu}_{i}\in\mathbb{R}^{d}$, is learned for each node $i\in\mathcal{V}$. For the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$, the latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ is a low-dimensional representation of $\bm{Y}_{i}\in\mathbb{R}^{n}$ at node $i$.
Here, the distribution of interest is $P(\bm{Y}_{i})$, which may be complicated. Instead of directly imposing a parametric form on the marginal $P(\bm{Y}_{i})$, our framework introduces an auxiliary latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ to model $P(\bm{Y}_{i})$ as a mixture:
$P(\bm{Y}_{i})=\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}.$
While the flexibility of our framework allows $\bm{Y}_{i}$ to be a vector of multivariate features, we regard $\bm{Y}_{i}\in\mathbb{R}^{n}$ as a univariate series of length $n$. Two remarks follow.
Remark 1.
The Gaussian assumption is placed on $P(\bm{Y}_{i}|\bm{Z}_{i})$, not on the marginal $P(\bm{Y}_{i})$. The induced distribution of $P(\bm{Y}_{i})$ can be highly non-Gaussian and depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$.
Remark 2.
Our framework assumes that $Y_{t,i}$ is independent in time conditional on $\bm{Z}_{i}$. Here, $\bm{Z}_{i}$ serves as our low-dimensional representation that captures temporal relationships while retaining a flexible marginal for $P(\bm{Y}_{i})$. In particular, the covariance of $\bm{Y}_{i}$ is
$\mathrm{Cov}(\bm{Y}_{i})$ $=\mathbb{E}\big[\mathrm{Cov}(\bm{Y}_{i}|\bm{Z}_{i})\big]+\mathrm{Cov}\big[\mathbb{E}(\bm{Y}_{i}|\bm{Z}_{i})\big]=\bm{I}_{n}+\mathrm{Cov}\big[\bm{h}_{\bm{\phi}}(\bm{Z}_{i})\big]$
which depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$, thereby allowing for marginal temporal dependence.
For flexibility, the mean of $P(\bm{Y}_{i}|\bm{Z}_{i})$ is modeled with neural networks, which have universal approximation properties (hornik1991approximation; yarotsky2017error). Specifically, $\bm{h}_{\bm{\phi}}(\bm{\cdot})$ is taken as a three-layer rectified linear unit (ReLU) neural network:
where $\bm{W}_{1}\in\mathbb{R}^{h_{1}\times d},\bm{W}_{2}\in\mathbb{R}^{h_{2}\times h_{1}}$, and $\bm{W}_{3}\in\mathbb{R}^{n\times h_{2}}$ are weights and $\bm{b}_{1}\in\mathbb{R}^{h_{1}\times 1},\bm{b}_{2}\in\mathbb{R}^{h_{2}\times 1}$, and $\bm{b}_{3}\in\mathbb{R}^{n\times 1}$ are bias parameters. Here, $\mathrm{ReLU}(x)=\max(0,x)$ is the element-wise ReLU activation function at $x$, and model parameters are collected into $\bm{\phi}$. While other choices of depth, activation functions, and architectures (lecun2015deep) can be used, ReLU neural networks are commonly employed and have been extensively studied.
@@
Figure 1 overviews our framework. The shaded circles in the top layer depict the series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$, and the dashed circle in the bottom layer depicts the latent $\{\bm{Z}_{i}\}_{i\in\mathcal{V}}$. The series are produced from the latent variables in a bottom-up manner. The decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with neural network parameter $\bm{\phi}$ is shared across all nodes, while the means $\bm{h}_{\bm{\phi}}(\bm{Z}_{i})$ differ by nodes. Intuitively, the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ helps learn the time series representation $\bm{Z}_{i}$ in a top-down manner.
$\bm{Z}_{1}$$\bm{Y}_{1}$$\mathcal{N}(\bm{\mu}_{1},\bm{I}_{d})$$\bm{Z}_{2}$$\bm{Y}_{2}$$\mathcal{N}(\bm{\mu}_{2},\bm{I}_{d})$$\bm{Z}_{3}$$\bm{Y}_{3}$$\mathcal{N}(\bm{\mu}_{3},\bm{I}_{d})$$\bm{Z}_{4}$$\bm{Y}_{4}$$\mathcal{N}(\bm{\mu}_{4},\bm{I}_{d})$$\bm{Z}_{5}$$\bm{Y}_{5}$$\mathcal{N}(\bm{\mu}_{5},\bm{I}_{d})$$P_{\bm{\phi}}(\bm{Y}_{1}|\bm{Z}_{1})$$P_{\bm{\phi}}(\bm{Y}_{2}|\bm{Z}_{2})$$P_{\bm{\phi}}(\bm{Y}_{3}|\bm{Z}_{3})$$P_{\bm{\phi}}(\bm{Y}_{4}|\bm{Z}_{4})$$P_{\bm{\phi}}(\bm{Y}_{5}|\bm{Z}_{5})$
@@
Figure 1: An illustration of our prior distributions and a decoder on a graph with five nodes.
@@ 3 Learning and Inference
$\ell(\bm{\phi},\bm{\mu}_{i})=\ln\big(P(\bm{Y}_{i};\bm{\phi},\bm{\mu}_{i})\big)=\ln\left(\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}\right).$
To facilitate clustering of the prior parameters, $\hat{\bm{\mu}}$, the penalized optimization
$\hat{\bm{\phi}},\hat{\bm{\mu}}=\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\left[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\mu}_{i}-\bm{\mu}_{j}\|_{2}\right]$ (1)
@@
To optimize (1) (which involves latent variables), we first manipulate the objective function. Introduce the slack variables $\bm{\nu}_{i,j}\in\mathbb{R}^{d}$ and recast the constrained optimization as
$\hat{\bm{\phi}},\hat{\bm{\mu}}=\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\left[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}\right],\qquad\text{subject to}\qquad\bm{\mu}_{i}-\bm{\mu}_{j}=\bm{\nu}_{i,j}.$ (2)
The augmented Lagrangian is
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{\rho})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+\sum_{(i,j)\in\mathcal{E}}\bm{\rho}_{i,j}^{\top}(\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j})+\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}\rVert_{2}^{2},$
where $\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ is the Lagrange multiplier for each $(i,j)\in\mathcal{E}$ and $\gamma>0$ is an additional penalty parameter for the augmentation term. Let $\bm{w}_{i,j}=\gamma^{-1}\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ denote a scaled dual variable. The augmented Lagrangian can be equivalently expressed as
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{w})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\left(\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}\rVert_{2}^{2}-\lVert\bm{w}_{i,j}\rVert_{2}^{2}\right).$
An ADMM procedure (boyd2011distributed) recursively solves our optimization:
$\bm{\phi}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\phi}}\left[-L(\bm{\phi},\bm{\mu}^{(a)})\right],$ (3)
$\bm{\mu}_{i}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\mu}_{i}}\left[-L(\bm{\phi}^{(a+1)},\bm{\mu}_{i})+\frac{\gamma}{2}\sum_{j\in\mathcal{B}(i)}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}^{(a)}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\right],\qquad\forall\ i\in\mathcal{V},$ (4)
$\bm{\nu}_{i,j}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\nu}_{i,j}}\left[\lambda\|\bm{\nu}_{i,j}\|_{2}+\frac{\gamma}{2}\lVert\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\right],\qquad\forall\ (i,j)\in\mathcal{E},$ (5)
$\bm{w}_{i,j}^{(a+1)}$ $=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\qquad\forall\ (i,j)\in\mathcal{E}.$ (6)
@@
$\bm{\mu}_{i}=\left(1+\gamma|\mathcal{B}(i)|\right)^{-1}\left[\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{j}+\bm{\nu}_{i,j}-\bm{w}_{i,j})\right].$ (7)
@@
The parameter $\bm{\phi}$ is updated via back-propagation. The update in (5) is equivalent to solving a group LASSO problem (yuan2006model; alaiz2013group). One can iteratively update $\bm{\nu}_{i,j}$ for each edge $(i,j)\in\mathcal{E}$ via

### v2

# Decoder-only Clustering in Graphs with Dynamic Attributes

@@ Abstract
This manuscript studies nodal clustering in graphs having a time series at each node. The proposed framework includes priors for low-dimensional representations and a decoder that bridges latent representations with time series. Addressing the limitation that the evolution of nodal attributes is often overlooked, temporal and structural patterns are fused into low-dimensional representations to facilitate clustering. Parameters are learned via maximum approximate likelihood, with a graph-fused LASSO regularization imposed on prior parameters. The optimization problem is solved via alternating direction method of multipliers; Langevin dynamics are employed for posterior inference. Simulation studies on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a word co-occurrence network demonstrate the effectiveness of the proposed clustering method.
ADMM, LASSO, Latent Space Models, Representation Learning, Time Series.
@@ 1 Introduction
Networks are used to describe relational phenomena, where entities are denoted by nodes and relations are delineated by edges. Such structures naturally arise in various domains, including social interactions (Snijders, 2011), neurology (Yang et al., 2022), and finance (Jackson and Pernoud, 2021). A fundamental task in network analysis involves partitioning nodes into disjoint sets, where nodes within a set are more connected than nodes in disjoint sets. Such clusters of nodes often provide valuable scientific insights (see Shai et al. (2021) for a collection of case studies). The majority of clustering methods focus on structural patterns (Malliaros and Vazirgiannis, 2013; Abbe, 2018; Macgregor, 2023). However, many networks possess dynamic nodal attributes, and the development of clustering attributed networks is still in its infancy. This paper demonstrates how to take advantage of nodal time series, providing clustering methods that account for both network structure and attribute dynamics. Section 5 applies the methods to networks with time-varying attributes, including identifying climate regions in California. Both proximities and monthly temperatures are used to develop a realistic clustering of the counties.
@@
Recent advances such as vision–language models highlight the importance of integrating heterogeneous sources of information in a coherent manner. Similarly, clustering nodes of attributed networks demands effective combination of graphical structures and nodal information. Methods have been previously proposed for clustering networks with static attributes. (Binkiewicz et al., 2017) perform covariate-assisted spectral clustering by augmenting the graph Laplacian with pairwise similarities of nodal covariates. (Shen et al., 2024) propose a Bayesian stochastic block model where nodal information is incorporated via a covariate-dependent random partition prior. (Hu and Wang, 2024) leverage individual and neighboring covariates with a node-specific weight matrix to improve clustering. Recently, the graph-fused LASSO regularizations in (Wang et al., 2016), (Chen et al., 2023), and (Madrid Padilla and Chen, 2023) have shown promising results, producing smoothed signals between neighboring nodes, while preserving boundaries between regions having distinct signals. Following (Hallac et al., 2015) and (Yu et al., 2025), our framework adapts the graph-fused LASSO regularization to achieve filtering behavior for clustering.
@@
In many analyses, nodal attributes progress in time, while the network structure remains fixed. Nodal time series can provide valuable clustering information, especially in cases with a non-informative graph such as a grid. Some methods for clustering nodes overlook the evolution of nodal features, flattening temporal patterns into static attributes. An intuitive solution introduces a latent variable at each node. By compressing each nodal series into a low-dimensional latent representation, temporal dynamics can be summarized and then integrated with the structural patterns to facilitate downstream inference. In our approach, latent representations are modeled by node-specific priors, and the latent space is regularized by a graph-fused LASSO penalty to incorporate network information. The resulting low-dimensional representations preserve both graphical structures and attribute dynamics, promoting effective nodal clustering.
@@
Latent space models naturally bridge low-dimensional representations with observed attributes. (Handcock et al., 2007) and (Sewell and Chen, 2015) model nodes in a Euclidean space, by assuming that edge probabilities are greater for node pairs that are closer in the latent space. (Wang et al., 2023) embed both nodes and their attributes in a shared latent space and develop a variational Bayesian EM algorithm to estimate parameters. (Zhu et al., 2023) construct a latent space model for both attractive and repulsive nodes, thereby disentangling network polarization dynamics. In particular, our framework models nodal time series, conditioning on their low-dimensional representations. Unlike a variational auto-encoder (Kingma and Welling, 2014; Kipf and Welling, 2016), which encourages the approximate posterior to align with a zero-mean Gaussian prior, we focus on learning the Gaussian prior means to facilitate clustering. With graph-fused LASSO regularization imposed on the differences between the prior parameters over edges, nodes having similar evolving patterns are encouraged to have similar prior parameters in Euclidean spaces. Consequently, the $k$-means method that clusters via centroids is well-suited in the resulting low-dimensional latent space.
@@
  • The associated optimization problem is solved via alternating direction method of multipliers (ADMM) with posterior inference. Through simulations on block and grid graphs with autoregressive dynamics, and applications to California county temperatures and a word co-occurrence network, the effectiveness of the proposed framework is demonstrated.
The rest of the manuscript is organized as follows. Section 2 describes our proposed framework. Section 3 presents an objective function with a graph-fused LASSO regularization and develops the ADMM procedure to solve the optimization problem. Section 4 discusses nodal clustering and model selection. Section 5 illustrates the proposed method through simulations and real-data applications. Section 6 discusses limitations and ideas for future work.
@@ 2 Decoder-only Latent Space Models
2.1 Model Specification
We consider an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are the respective node and edge sets. For each node $i\in\mathcal{V}$, let $\{Y_{t,i}\}_{t=1}^{n}$ denote a sequence of univariate random variables. The node $i$ series is represented as $\bm{Y}_{i}\coloneqq(Y_{1,i},Y_{2,i},\dots,Y_{n,i})^{\top}$. For each $\bm{Y}_{i}\in\mathbb{R}^{n}$, we assume existence of a latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ such that $\bm{Y}_{i}$ is generated from $\bm{Z}_{i}$ with the following time series decoder:
@@ 2.1 Model Specification
where $\bm{h}_{\bm{\phi}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$ is a neural network parameterized by $\bm{\phi}$ and is elaborated in Section 2.2. Here, $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ is the conditional probability density function of $\bm{Y}_{i}$ given $\bm{Z}_{i}$. We impose a prior distribution on the latent $\bm{Z}_{i}$ via
@@
The prior parameter, $\bm{\mu}_{i}\in\mathbb{R}^{d}$, is learned for each node. For the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$, the latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ is a low-dimensional representation of $\bm{Y}_{i}\in\mathbb{R}^{n}$ at node $i$.
@@
Figure 1 overviews our framework. The shaded circles in the top layer depict the series $\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$, and the dashed circles in the bottom layer depict the latent $\{\bm{Z}_{i}\}_{i\in\mathcal{V}}$. The series are generated from the latent variables in a bottom-up manner. The decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ with neural network parameter $\bm{\phi}$ is shared across all nodes, while the means $\bm{h}_{\bm{\phi}}(\bm{Z}_{i})$ differ by nodes. Intuitively, the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ helps learn the time series representation $\bm{Z}_{i}\in\mathbb{R}^{d}$ in a top-down manner. The simplicity of the proposed framework is noted, without the need for an encoder.
[…]
@@
Figure 1: An illustration of our prior distributions and a decoder on a graph with five nodes and two clusters.
2.2 Neural Decoder and Temporal Dependence
The distribution of interest is the marginal $P(\bm{Y}_{i})$, which is often complicated. Instead of imposing a parametric form directly on the marginal $P(\bm{Y}_{i})$, our framework introduces an auxiliary latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ to model $P(\bm{Y}_{i})$ from a mixture:
$P(\bm{Y}_{i})=\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}$ (1)
with the following remarks.
Remark 2.1.
The Gaussian assumption is placed on the conditional $P(\bm{Y}_{i}|\bm{Z}_{i})$, instead of the marginal $P(\bm{Y}_{i})$. The induced distribution of $P(\bm{Y}_{i})$ can be highly non-Gaussian and depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$.
Remark 2.2.
Our framework assumes that $Y_{t,i}$ is independent in time $t$ conditional on $\bm{Z}_{i}$. Here, $\bm{Z}_{i}\in\mathbb{R}^{d}$ serves as a low-dimensional representation that captures temporal relationships while retaining a flexible marginal for $P(\bm{Y}_{i})$. In particular, the covariance of $\bm{Y}_{i}\in\mathbb{R}^{n}$ is
$\mathrm{Cov}(\bm{Y}_{i})$ $=\mathbb{E}\big[\mathrm{Cov}(\bm{Y}_{i}|\bm{Z}_{i})\big]+\mathrm{Cov}\big[\mathbb{E}(\bm{Y}_{i}|\bm{Z}_{i})\big]$
$=\bm{I}_{n}+\mathrm{Cov}\big[\bm{h}_{\bm{\phi}}(\bm{Z}_{i})\big]$
which depends on $\bm{h}_{\bm{\phi}}(\bm{\cdot})$, thereby allowing for potentially complex marginal temporal dependence.
Remark 2.3.
The negative log-likelihood of (1) that involves latent $\bm{Z}_{i}$ is lower bounded (Mattei and Frellsen, 2018):
$\ln\big(P(\bm{Y}_{i})\big)\geq\frac{n}{2}\ln(2\pi),$
rendering associated minimization problems meaningful.
For flexibility of the marginal $P(\bm{Y}_{i})$, the mean $\bm{h}_{\bm{\phi}}(\bm{\cdot})$ in the decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$ is parameterized by a three-layer ReLU neural network, which have universal approximation properties (Hornik et al., 1989; Hornik, 1991; Yarotsky, 2017). This parameterization is particularly chosen for representation learning to capture temporal patterns, yielding node-level representations tailored for clustering and graph-based regularization. In contrast, recurrent-based neural networks are typically designed to model conditional dynamics and are optimized for sequential forecasting tasks, making the latent representations less amenable to node-level abstraction for graph related tasks. Furthermore, recurrent-based neural networks are often computationally more demanding, jeopardizing the optimization procedure elaborated in Section 3. While other architectures can be used, ReLU neural networks have been extensively studied and is well suited for the proposed clustering framework.
@@ 3 Learning and Inference
$\ell(\bm{\phi},\bm{\mu}_{i})=\ln\left(\int P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})P_{\bm{\mu}_{i}}(\bm{Z}_{i})d\bm{Z}_{i}\right).$
To facilitate clustering of the prior parameters, the penalized optimization
$\hat{\bm{\phi}},\hat{\bm{\mu}}=\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\bigg[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\mu}_{i}-\bm{\mu}_{j}\|_{2}\bigg]$ (2)
@@
To optimize (2) which involves latent variables, we first manipulate the objective function. Introduce the slack variables $\bm{\nu}_{i,j}\in\mathbb{R}^{d}$ and recast the constrained optimization as
$\begin{split}\hat{\bm{\phi}},\hat{\bm{\mu}}=&\operatorname*{arg\,min}_{\bm{\phi},\bm{\mu}}\left[-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}\right],\\ &\text{subject to}\quad\bm{\mu}_{i}-\bm{\mu}_{j}=\bm{\nu}_{i,j}.\end{split}$ (3)
The augmented Lagrangian is formulated as
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{\rho})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+$
$\sum_{(i,j)\in\mathcal{E}}\bm{\rho}_{i,j}^{\top}(\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j})+\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}\rVert_{2}^{2},$
where $\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ is the Lagrange multiplier for each $(i,j)\in\mathcal{E}$ and $\gamma>0$ is an additional penalty parameter for the augmentation term. Let $\bm{w}_{i,j}=\gamma^{-1}\bm{\rho}_{i,j}\in\mathbb{R}^{d}$ denote the scaled dual variable. The augmented Lagrangian can be equivalently formulated as
$\mathcal{L}(\bm{\phi},\bm{\mu},\bm{\nu},\bm{w})=-L(\bm{\phi},\bm{\mu})+\lambda\sum_{(i,j)\in\mathcal{E}}\|\bm{\nu}_{i,j}\|_{2}+$
$\frac{\gamma}{2}\sum_{(i,j)\in\mathcal{E}}\left(\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}+\bm{w}_{i,j}\rVert_{2}^{2}-\lVert\bm{w}_{i,j}\rVert_{2}^{2}\right).$
An ADMM procedure (Boyd et al., 2011; Wang et al., 2019) recursively solves our optimization:
$\bm{\phi}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\phi}}\bigg[-L(\bm{\phi},\bm{\mu}^{(a)})\bigg],$
$\bm{\mu}_{i}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\mu}_{i}}\bigg[-L(\bm{\phi}^{(a+1)},\bm{\mu}_{i})+$ (4)
$\frac{\gamma}{2}$ $\sum_{j\in\mathcal{B}(i)}\lVert\bm{\mu}_{i}-\bm{\mu}_{j}-\bm{\nu}_{i,j}^{(a)}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\bigg],\,\,\forall\ i\in\mathcal{V},$
$\bm{\nu}_{i,j}^{(a+1)}$ $=\operatorname*{arg\,min}_{\bm{\nu}_{i,j}}\bigg[\lambda\|\bm{\nu}_{i,j}\|_{2}+$ (5)
$\frac{\gamma}{2}$ $\lVert\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}+\bm{w}_{i,j}^{(a)}\rVert_{2}^{2}\bigg],\,\,\forall\ (i,j)\in\mathcal{E},$
$\bm{w}_{i,j}^{(a+1)}$ $=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\,\,\forall\ (i,j)\in\mathcal{E}.$
@@
$\bm{\mu}_{i}=\frac{1}{(1+\gamma|\mathcal{B}(i)|)}\bigg[\mathbb{E}(\bm{Z}_{i}|\bm{Y}_{i})+\gamma\sum_{j\in\mathcal{B}(i)}(\bm{\mu}_{j}+\bm{\nu}_{i,j}-\bm{w}_{i,j})\bigg].$ (6)
@@
The parameter $\bm{\phi}$ is updated via back-propagation. The update in (5) is equivalent to solving a group LASSO problem (Yuan and Lin, 2006; Vert and Bleakley, 2010; Alaíz et al., 2013). One can update $\bm{\nu}_{i,j}$ for each edge $(i,j)\in\mathcal{E}$ via